OpenStudy (narissa):

evaluate the expression -3x ÷ y for each given x and y value. a) if x =10 and y=-3 b) if x =-10 and y =3 c) if x =-10 and y =-3

OpenStudy (narissa):

@HossamHoussien

OpenStudy (narissa):

@HossamHoussien

OpenStudy (narissa):

@Willie579

OpenStudy (hossamhoussien):

a. \[(10,-3)\rightarrow \frac{ -3(10) }{ -3 }=10\] b. \[(-10,3)\rightarrow \frac{ -3(-10) }{ 3 }=10\] c. \[(-10,-3)\rightarrow \frac{ -3(-10) }{ -3 }=-10\]

OpenStudy (mathmale):

@HossamHoussien: Please give at least one example and then engage your student in the learning process. Do NOT do the work of others for them and then give away answers, as you have done here. Thank you.